# Cislunar Space Situational Awareness Constellation Design and Planning with Facility Location Problem

### v3

@@ Abstract
[…]
@@
[…]
@@ 1.1 Overview of Existing Approaches for Cislunar Space Situational Awareness Problems
In an LP-based approach, an LP model, commonly involving binary or integer variables, is constructed and solved [30, 21, 31, 20]. Nonlinearities surrounding either the constellation design or sensor-tasking, are precomputed to form objectives and/or constraints to the model. Mixed-integer LP (MILP) and binary LP (BLP) are well-suited for encoding combinatorial decisions, and a careful formulation can lead to problem formulations that can be efficiently solved by specialized algorithms such as branch and bound (BnB) [32, 33, 34, 35], Lagrangian method (LM) [36, 37, 38], or column generation [39, 40, 41]. Additionally, LP algorithms provide an optimality gap, which corresponds to the worst-case degradation of a prematurely terminated suboptimal solution from the predicted optimal solution. Such a gap is not readily available in NLP-based approaches such as EA, MCTS, or ML. Notable challenges surrounding LP-based approaches include the need for a reasonable discretization scheme to approximate the continuous decision space in observer placement and sensor tasking, and devising an appropriate formulation that is tractable.
@@ 1.2 Problem Definition and Proposed Approach
While generic BnB algorithms are known to adequately tackle classical $p$-median problems of moderate size, we also propose a custom LM to compute an approximate solution rapidly. LM consists of iteratively solving for an idealized optimal solution to a relaxed problem and a heuristics-based feasible solution, each time updating Lagrange multipliers used to construct the relaxed problem. In this paper, an LM with a purely analytical solution to the relaxed problem is derived by decoupling the observer-wise sensor-tasking decision. Custom heuristics for the feasible solution are designed to leverage the orbital structure of candidate observer locations. We note that the large number of variables, in the order of $10^{6}$, of the TE-$p$-MP prohibits the use of EAs. An NLP-based approach for tackling both constellation design and sensor-tasking problems could consist, for example, of formulating a black-box constellation design problem in the outer loop, with sensor-tasking handled in the inner loop by a pre-trained neural network (NN), or a policy chosen from a predefined set by the outer loop. We however opt for the proposed LP-based approach in favor of its ease in constructing a tractable formulation that encodes decisions surrounding the observation schedule. With an NLP-based approach, a policy-based approach would result in a limited performance due to the inherently enforced structure by the pre-defined set of policies, while parameterizing the policy to be optimized by an outer loop is non-trivial. Recently, Patel et al. [20] also proposed a framework for concurrent CSSA constellation design and sensor-tasking. Their approach is a hybrid between NLP and LP, where an outer-layer NLP optimizes the phasing of satellites along pre-defined orbits, and an inner-layer LP assigns discrete targets to observers. This work differs from [20] in the following ways: our formulation optimizes not only for the phasing but also the orbits of observers from a discrete set of options; we also consider target observation by considering a sensor FOV, rather than assuming a one-to-one assignment between observers and targets; finally, our formulation is a single-layer MILP that does not rely on a black-box NLP solver. Through the proposed TE-$p$-MP together with the LM, our approach can produce a feasible candidate CSSA constellation design together with a pointing direction schedule for each observer that provides adequate coverage to a given set of demand and observer parameters.
@@ 2.1.1 $p$-Median Problem
$Y_{j}$ =$=\begin{cases}1&\text{facility location $j$ is used}\\ 0&\text{otherwise}\end{cases}$ (1)
$X_{ij}$ =$=\begin{cases}1&\text{client $i$ is allocated to facility location $j$}\\ 0&\text{otherwise}\end{cases}$ (2)
@@ 2.2 Lagrangian Method for Integer Linear Program
$\max_{\boldsymbol{z}}$ $\quad\boldsymbol{f}^{T}\boldsymbol{z}$ (4a)
@@
$\max_{\boldsymbol{z}}$ $\quad\boldsymbol{f}^{T}\boldsymbol{z}-\boldsymbol{\lambda}^{T}\left(\boldsymbol{b}_{c}-\boldsymbol{A}_{c}\boldsymbol{z}\right)$ (5)
@@ 2.3.1 Equations of Motion
[…]
@@ 2.3.2 Libration Point Orbit
Multiple known families of LPOs exist in the vicinity of L1 and L2; among the members of each family, a discrete subset of LPOs have periods approximately equal to an integer ratio of the synodic period $T_{\mathrm{syn}}$, the duration between two consecutive Sun-Earth-Moon alignment. We denote an $M$:$N$ LPO as an orbit with period $T=(N/M)T_{\mathrm{syn}}$. In this work, we focus on a subset of such $M$:$N$ resonant LPOs as part of the candidate observer orbits, following previous works reporting their favorable illumination conditions for CSSA applications [2, 13]. We note that this choice is reminiscent of opting for designing Earth observation constellations with repeating Low-Earth orbits that exhibit repeating ground tracks [57]. An actual spacecraft deployed in a cislunar constellation will be subject to full-ephemeris dynamics subject to various perturbations. Furthermore, due to the Moon’s and Earth’s eccentricities, $T_{\mathrm{syn}}$ is not constant. Thus, in reality, the $M$:$N$ resonance is only an approximation. Nevertheless, we assume the spacecraft will be actively controlled via occasional station-keeping maneuvers to remain in the approximate vicinity of the quasi-periodic LPO that geometrically resembles the CR3BP LPO. In this work, we thus assume the illumination conditions after $M$ revolutions for a given $M$:$N$ resonant LPO are perfectly repeating. The interested reader is pointed to Klonowski et al. [25] for an analysis of performance variation due to this drift.
@@
This work considers the synodic-resonant LPOs in the L1/L2 Lyapunov, Southern and Northern L2 Halo, Southern and Northern Butterfly, Distant Prograde Orbit (DPO), and Distant Retrograde Orbit (DRO) families, with synodic resonances 1:1, 3:2, 2:1, 9:4, 5:2, 3:1, 4:1, and 9:2. This selection results in 40 LPOs, as shown in Figure 1. The parameters of these LPOs are given in Appendix 7.1. Each LPO is constructed via a fixed-time variant of the single-shooting differential correction approach [58], using the LPO entry in the JPL Three-Body Periodic Orbit Catalog [^4] https://ssd.jpl.nasa.gov/tools/periodic_orbits.html with the closest period to the sought resonant period as initial guess. The choice of which resonant LPOs to include as candidate orbits matters if the relative positions of member satellites within a constellation must be repeated over a $N_{\rm syn}$-multiple of the synodic month, where $N_{\rm syn}$ is the smallest integer such that the product $N_{\rm syn}M/N$ is an integer for all considered $M$:$N$ ratios. This translates to whether the sought observer location and pointing schedule are only required to be transient, or if a repeating, steady-state solution is sought. With our choice of resonant LPOs, the constellation returns exactly to its initial configuration every 4 synodic months. We note that the TE-$p$-MP formulation developed in later Sections does not explicitly depend on the $N_{\rm syn}$-month repeating pattern of candidate LPOs, and thus one may instead choose to use an arbitrary set of candidate observer LPOs. However, the number of discretized time steps necessitated to solve a CSSA constellation and sensor-tasking over an extended duration, for example, lasting multiple years, can be reduced by several folds.
@@ 3.2 Dynamic Demand: Low-Energy Transfer Transit Window
[…]
@@
Figure 4: Sparsity of demand matrix $D\in\mathbb{B}^{\ell\times q}$ for LET transit window over 120 time steps corresponding to $4$ synodic months
@@ 3.3.3 Visibility boolean
$\bar{m}^{b}(\boldsymbol{r}_{j}(t),\boldsymbol{r}_{k}(t),i(t))$ =$=\begin{cases}0&\cos^{-1}\left(\boldsymbol{l}_{jM}(t)\cdot\boldsymbol{l}_{jk}(t)\right)<\phi_{\mathrm{Moon}}^{\mathrm{crit}}(\boldsymbol{r}_{j})\\ 0&\cos^{-1}\left(\boldsymbol{l}_{jE}(t)\cdot\boldsymbol{l}_{jk}(t)\right)<\phi_{\mathrm{Earth}}^{\mathrm{crit}}(\boldsymbol{r}_{j})\\ 0&\boldsymbol{r}_{k}(t)\notin\operatorname{FOVCone}_{i}\left(\boldsymbol{r}_{j}(t)\right)\\ 0&\bar{m}_{\mathrm{target}}>\bar{m}_{\mathrm{crit}}\\ 1&\text{otherwise}\end{cases}$ (19)
@@ 4.1 Design Variables
$X_{ijt}$ =$=\begin{cases}1&\text{observer in location $j$ observes along direction $i$ at time $t$}\\ 0&\text{otherwise}\end{cases}$ (20)
@@ 4.2 Problem Formulation
$\max_{X,Y}\quad$ $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\left(\max_{\begin{subarray}{c}i=1,\ldots,m,\\ j=1,\ldots,n\end{subarray}}M^{b}_{ijtk}X_{ijt}\right)-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}$ (22a)
@@
$\max_{X,Y,\theta}\quad$ $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\theta_{tk}-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}$ (24a)
@@
(22b) -(22e)
where constraint (24b) ensures this modified problem (24) has the same optimal solution as the original problem (22). Problem (24) is a standard mixed-integer linear program (MILP), which is amenable to standard BnB solvers such as Gurobi [65] or CPLEX [66]. While the slack variable $\theta_{tk}$ is a continuous variable in $[0,1]$, the optimal solution to (24) has $\theta_{tk}=1$ for time $t$ and target $k$ if at least one observer is viewing $k$ at this time, and $\theta_{tk}=0$ if $\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}=0$. Thus, this slack variable may be interpreted as a binary variable, such that
$\theta_{tk}=\begin{cases}1&\text{target $k$ is observed by at least one observer at time $t$}\\ 0&\text{otherwise}\end{cases}$ (25)
@@ 5.1 Upper Bound Solution via Lagrangian Relaxation
$\max_{X,Y,\theta}$ $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\theta_{tk}-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}-\sum_{j=1}^{n}\sum_{t=1}^{\ell}\lambda_{jt}\left(\sum_{i=1}^{m}X_{ijt}-1\right)-\sum_{t=1}^{\ell}\sum_{k=1}^{q}\eta_{tk}\left(\theta_{tk}-\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}\right)$ (27)
s.t. (22c) -(22e), (24c)
@@
$\max_{X,Y,\theta}\quad\sum_{t=1}^{\ell}\sum_{k=1}^{q}\left(1-\eta_{tk}\right)\theta_{tk}+\sum_{i=1}^{m}\sum_{j=1}^{n}\sum_{t=1}^{\ell}\left(\sum_{k=1}^{q}\eta_{tk}M^{b}_{ijtk}-\lambda_{jt}\right)X_{ijt}-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}+\sum_{j=1}^{n}\sum_{t=1}^{\ell}\lambda_{jt}$ (28)
@@ 5.1.1 Decomposition into Subproblems
$\max_{X,Y_{j},\theta}\quad$ $\sum_{i=1}^{m}\sum_{t=1}^{\ell}\left(\sum_{k=1}^{q}\eta_{tk}M^{b}_{ijtk}-\lambda_{jt}\right)X_{ijt}-\dfrac{1}{\ell}\left(f_{j}Y_{j}\right)+\sum_{t=1}^{\ell}\lambda_{jt}$ (30a)
@@ 5.1.2 Construction of Lagrangian Relaxed Solution
1:$\lambda_{jt}$ $\forall j,t$, $\eta_{tk}$ $\forall t,k$
2:$\bar{Y},\bar{X}\leftarrow\boldsymbol{0}_{n\times 1},\boldsymbol{0}_{m\times n\times\ell}$triangleright Initialize
@@
8:$\pi\leftarrow\verb|argsortDescend|(Z_{\rm relax}^{1},\ldots,Z_{\rm relax}^{n})$triangleright Sort in descending order according to (33)
@@
10:     $\bar{Y}_{j}\leftarrow 1$triangleright Use facility location $j$
11:     $\bar{X}_{ijt}\leftarrow\bar{X}_{ijt}^{j}$triangleright Assign targets according to decision made for subproblem $j$

### v4

@@ Abstract
FOV  = field of view,  $\text{\,}\mathrm{d}\mathrm{e}\mathrm{g}$
@@
$\phi$  = phase angle,  $\text{\,}\mathrm{d}\mathrm{e}\mathrm{g}$
@@ 1.1 Overview of Existing Approaches for Cislunar Space Situational Awareness Problems
In an LP-based approach, an LP model, commonly involving binary or integer variables, is constructed and solved [30, 21, 31, 20]. Nonlinearities surrounding either the constellation design or sensor-tasking, are precomputed to form objectives and/or constraints to the model. Mixed-integer LP (MILP) and binary LP (BLP) are well-suited for encoding combinatorial decisions, and a careful formulation can lead to problem formulations that can be efficiently solved by specialized algorithms such as branch and bound (B&B) [32, 33, 34, 35], Lagrangian method (LM) [36, 37, 38], or column generation [39, 40, 41]. Additionally, LP algorithms provide an optimality gap, which corresponds to the worst-case degradation of a prematurely terminated suboptimal solution from the predicted optimal solution. Such a gap is not readily available in NLP-based approaches such as EA, MCTS, or ML. Notable challenges surrounding LP-based approaches include the need for a reasonable discretization scheme to approximate the continuous decision space in observer placement and sensor tasking, and devising an appropriate formulation that is tractable.
@@ 1.2 Problem Definition and Proposed Approach
While generic B&B algorithms are known to adequately tackle classical $p$-median problems of moderate size, we also propose a custom LM to compute an approximate solution rapidly. LM consists of iteratively solving for an idealized optimal solution to a relaxed problem and a heuristics-based feasible solution, each time updating Lagrange multipliers used to construct the relaxed problem. In this paper, an LM with a purely analytical solution to the relaxed problem is derived by decoupling the observer-wise sensor-tasking decision. Custom heuristics for the feasible solution are designed to leverage the orbital structure of candidate observer locations. We note that the large number of variables, in the order of $10^{6}$, of the TE-$p$-MP prohibits the use of EAs. An NLP-based approach for tackling both constellation design and sensor-tasking problems could consist, for example, of formulating a black-box constellation design problem in the outer loop, with sensor-tasking handled in the inner loop by a pre-trained neural network (NN), or a policy chosen from a predefined set by the outer loop. We however opt for the proposed LP-based approach in favor of its ease in constructing a tractable formulation that encodes decisions surrounding the observation schedule. With an NLP-based approach, a policy-based approach would result in a limited performance due to the inherently enforced structure by the pre-defined set of policies, while parameterizing the policy to be optimized by an outer loop is non-trivial. Recently, Patel et al. [20] also proposed a framework for concurrent CSSA constellation design and sensor-tasking. Their approach is a hybrid between NLP and LP, where an outer-layer NLP optimizes the phasing of satellites along pre-defined orbits, and an inner-layer LP assigns discrete targets to observers. This work differs from [20] in the following ways: our formulation optimizes not only for the phasing but also the orbits of observers from a discrete set of options; we also consider target observation by considering a sensor FOV, rather than assuming a one-to-one assignment between observers and targets; finally, our formulation is a single-layer MILP that does not rely on a black-box NLP solver. Through the proposed TE-$p$-MP together with the LM, our approach can produce a feasible candidate CSSA constellation design together with a pointing direction schedule for each observer that provides adequate coverage to a given set of demand and observer parameters.
@@ 2.1.1 $p$-Median Problem
$Y_{j}$ $=\begin{cases}1&\text{facility location $j$ is used}\\ 0&\text{otherwise}\end{cases}$ (1)
$X_{ij}$ $=\begin{cases}1&\text{client $i$ is allocated to facility location $j$}\\ 0&\text{otherwise}\end{cases}$ (2)
@@ 2.2 Lagrangian Method for Integer Linear Program
$\underset{\boldsymbol{z}}{\operatorname{maximize}}$ $\quad\boldsymbol{f}^{T}\boldsymbol{z}$ (4a)
@@
$\underset{\boldsymbol{z}}{\operatorname{maximize}}$ $\quad\boldsymbol{f}^{T}\boldsymbol{z}-\boldsymbol{\lambda}^{T}\left(\boldsymbol{b}_{c}-\boldsymbol{A}_{c}\boldsymbol{z}\right)$ (5)
@@ 2.3.1 Equations of Motion
Length unit $\mathrm{LU}$,  $\text{\,}\mathrm{k}\mathrm{m}$ 389703.2648292776
Time unit $\mathrm{TU}$,  $\text{\,}\mathrm{s}$ 382981.2891290545
Synodic period,  $\text{\,}\mathrm{d}\mathrm{a}\mathrm{y}$ 29.5
@@ 2.3.2 Libration Point Orbit
Multiple known families of LPOs exist in the vicinity of L1 and L2; among the members of each family, a discrete subset of LPOs have periods approximately equal to an integer ratio of the synodic period $T_{\mathrm{syn}}$, the duration between two consecutive Sun-Earth-Moon alignment. We denote an $M$:NN LPO as an orbit with period $T=(N/M)T_{\mathrm{syn}}$. In this work, we focus on a subset of such $M$:NN resonant LPOs as part of the candidate observer orbits, following previous works reporting their favorable illumination conditions for CSSA applications [2, 13]. We note that this choice is reminiscent of opting for designing Earth observation constellations with repeating Low-Earth orbits that exhibit repeating ground tracks [57]. An actual spacecraft deployed in a cislunar constellation will be subject to full-ephemeris dynamics subject to various perturbations. Furthermore, due to the Moon’s and Earth’s eccentricities, $T_{\mathrm{syn}}$ is not constant. Thus, in reality, the $M$:NN resonance is only an approximation. Nevertheless, we assume the spacecraft will be actively controlled via occasional station-keeping maneuvers to remain in the approximate vicinity of the quasi-periodic LPO that geometrically resembles the CR3BP LPO. In this work, we thus assume the illumination conditions after $M$ revolutions for a given $M$:NN resonant LPO are perfectly repeating. The interested reader is pointed to Klonowski et al. [25] for an analysis of performance variation due to this drift.
@@
This work considers the synodic-resonant LPOs in the L1/L2 Lyapunov, Southern and Northern L2 Halo, Southern and Northern Butterfly, Distant Prograde Orbit (DPO), and Distant Retrograde Orbit (DRO) families, with synodic resonances 1:1, 3:2, 2:1, 9:4, 5:2, 3:1, 4:1, and 9:2. This selection results in 40 LPOs, as shown in Figure 1. The parameters of these LPOs are given in Appendix 7.1. Each LPO is constructed via a fixed-time variant of the single-shooting differential correction approach [58], using the LPO entry in the JPL Three-Body Periodic Orbit Catalog [^4] https://ssd.jpl.nasa.gov/tools/periodic_orbits.html with the closest period to the sought resonant period as initial guess. The choice of which resonant LPOs to include as candidate orbits matters if the relative positions of member satellites within a constellation must be repeated over a $N_{\rm syn}$-multiple of the synodic month, where $N_{\rm syn}$ is the smallest integer such that the product $N_{\rm syn}M/N$ is an integer for all considered $M$:NN ratios. This translates to whether the sought observer location and pointing schedule are only required to be transient, or if a repeating, steady-state solution is sought. With our choice of resonant LPOs, the constellation returns exactly to its initial configuration every 4 synodic months. We note that the TE-$p$-MP formulation developed in later Sections does not explicitly depend on the $N_{\rm syn}$-month repeating pattern of candidate LPOs, and thus one may instead choose to use an arbitrary set of candidate observer LPOs. However, the number of discretized time steps necessitated to solve a CSSA constellation and sensor-tasking over an extended duration, for example, lasting multiple years, can be reduced by several folds.
@@ 3.2 Dynamic Demand: Low-Energy Transfer Transit Window
LETs are translunar trajectories that leverage the weak stability boundary of the Earth-Moon-Sun system to reduce the arrival specific energy with respect to the Moon at the cost of longer times of flight [61, 62, 63]. A sample set of LETs computed with the bi-circular restricted four-body model is shown in Figure 3 in the Earth-Moon rotating frame. The left pane in the Figure shows the entire trajectory from perigee to perilune at $100\text{\,}\mathrm{k}\mathrm{m}$ altitude; the center and right pane shows the portion of each LETs in the lunar vicinity. Due to their high apogee of around $1.5$ million km, LETs are particularly hard to monitor during the majority of their transfer, even with the existence of an in-space SSA network capable of observing areas such as the lunar SOI or the cone of shame. One strategic location to monitor LETs is the neck region around Earth-Moon L2, where trajectories pass through before entering the cislunar vicinity. The center and right pane in Figure 3 shows this behavior. We consider monitoring a volume corresponding to this neck region-based transit window, centered at L2 and spanning $20\,000\text{\,}\mathrm{k}\mathrm{m}$ along the $x$-axis and $100\,000\text{\,}\mathrm{k}\mathrm{m}$ along the $y$ and $z$-axes of the Earth-Moon rotating frame.
@@
Figure 4: Sparsity of demand matrix $D\in\mathbb{B}^{\ell\times q}$ for LET transit window over 120 time steps corresponding to 44 synodic months
@@ 3.3.3 Visibility boolean
[…]
@@ 4.1 Design Variables
[…]
@@ 4.2 Problem Formulation
$\underset{X,Y}{\operatorname{maximize}}\quad$ $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\left(\max_{\begin{subarray}{c}i=1,\ldots,m,\\ j=1,\ldots,n\end{subarray}}M^{b}_{ijtk}X_{ijt}\right)-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}$ (22a)
@@
$\underset{X,Y,\theta}{\operatorname{maximize}}\quad$ $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\theta_{tk}-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}$ (24a)
@@
(22b) $\sim$ (22e)
where constraint (24b) ensures this modified problem (24) has the same optimal solution as the original problem (22). Problem (24) is a standard mixed-integer linear program (MILP), which is amenable to standard B&B solvers such as Gurobi [65] or CPLEX [66]. While the slack variable $\theta_{tk}$ is a continuous variable in $[0,1]$, the optimal solution to (24) has $\theta_{tk}=1$ for time $t$ and target $k$ if at least one observer is viewing $k$ at this time, and $\theta_{tk}=0$ if $\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}=0$. Thus, this slack variable may be interpreted as a binary variable, such that
[…]
@@ 5.1 Upper Bound Solution via Lagrangian Relaxation
$\underset{X,Y,\theta}{\operatorname{maximize}}$ $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\theta_{tk}-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}-\sum_{j=1}^{n}\sum_{t=1}^{\ell}\lambda_{jt}\left(\sum_{i=1}^{m}X_{ijt}-1\right)-\sum_{t=1}^{\ell}\sum_{k=1}^{q}\eta_{tk}\left(\theta_{tk}-\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}\right)$ (27)
s.t. (22c) $\sim$ (22e), (24c)
@@
$\underset{X,Y,\theta}{\operatorname{maximize}}\quad\sum_{t=1}^{\ell}\sum_{k=1}^{q}\left(1-\eta_{tk}\right)\theta_{tk}+\sum_{i=1}^{m}\sum_{j=1}^{n}\sum_{t=1}^{\ell}\left(\sum_{k=1}^{q}\eta_{tk}M^{b}_{ijtk}-\lambda_{jt}\right)X_{ijt}-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}+\sum_{j=1}^{n}\sum_{t=1}^{\ell}\lambda_{jt}$ (28)
@@ 5.1.1 Decomposition into Subproblems
$\underset{X,Y_{j},\theta}{\operatorname{maximize}}\quad$ $\sum_{i=1}^{m}\sum_{t=1}^{\ell}\left(\sum_{k=1}^{q}\eta_{tk}M^{b}_{ijtk}-\lambda_{jt}\right)X_{ijt}-\dfrac{1}{\ell}\left(f_{j}Y_{j}\right)+\sum_{t=1}^{\ell}\lambda_{jt}$ (30a)
@@ 5.1.2 Construction of Lagrangian Relaxed Solution
[…]
@@
[…]
@@
10:  $\bar{Y}_{j}\leftarrow 1$ $\triangleright$ Use facility location $j$
11:  $\bar{X}_{ijt}\leftarrow\bar{X}_{ijt}^{j}$ $\triangleright$ Assign targets according to decision made for subproblem $j$